OpenStudy (666potatoes):

Rationalize the denominator: √5 / (√5 + 2) [2 is not part of the square root of 5]

OpenStudy (anonymous):

multiply both sides by \[\sqrt{5} - 2\]

OpenStudy (anonymous):

rachel is right, multiply by the reciprocal

OpenStudy (666potatoes):

Would it leave me with √25 - 2 ?

OpenStudy (666potatoes):

Or √25 - 2 / √25 - 4 ?

OpenStudy (anonymous):

the radicals cancel

OpenStudy (666potatoes):

Ah that's what I thought. So it is just √25 - 2 ? so it would be 3?

OpenStudy (anonymous):

\[25-2\sqrt{5}\]

OpenStudy (anonymous):

(You had to multiply root 5 by -2)

OpenStudy (anonymous):

how do you get 25-2root5...i got 5-2root5..

OpenStudy (666potatoes):

Zomg, a dispute.

OpenStudy (666potatoes):

I don't think the radicals cancel out when multiplying √5 by √5 in the numerator. :l

OpenStudy (666potatoes):

But then that means the denominator radicals wouldn't cancel out so weogheoiwghoiew.

OpenStudy (anonymous):

They do! if you multiply the same radicals together they drop the radical sign

OpenStudy (666potatoes):

Alright, I'm going with Rachel's answer. Haha.

OpenStudy (anonymous):

i agree with rachel, i just get the 25...cuz in the numerator i was getting root 5 times root5 -2. root 5 times root 5 =5 not 25

OpenStudy (anonymous):

just not*

OpenStudy (anonymous):

Oh right. sorry i messed up.

OpenStudy (anonymous):

haha its cool..i wrote that at first too

OpenStudy (anonymous):

all rightie! good luck with math, and finals if you have them!

OpenStudy (666potatoes):

So the correct answer is 5-2√5. Right?

OpenStudy (anonymous):

Yep!

OpenStudy (666potatoes):

Ty gaiz. c: